How do I work this? I am drawing a blank. Thank You for any help.
sin θ = - ½ , 180° < 0 < 270°, find tan θ

Answer by stanbon(75887)

sin θ = - ½ , 180° < 0 < 270°, find tan θ
-----
Theta is in the 3rd Quadrant where x and y are both negative.
---
Since sin = y/r, y = -1 and r = 2
---
solve for "x":
2^2 = (-1)^2 + x^2
---
x = -sqrt(3) (negative because theta is in the 3rd quadrant)
-------
Then tan(theta) = y/x = -1/(-sqrt(3)) = 1/sqrt(3) = (1/3)sqrt(3)
==================
Cheers,
Stan H.